Sol Summary- May 02nd

Crew 211 Sol 11 Summary Report 2-May-2019

Sol: 11

Summary Title: Sunset on Mars

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: The commander and vice-commander began the day with a nice morning EVA along the Hab Ridge Road. After the breakfast, the commander and vice-commander deployed their respective experiment during an engineering EVA while the rest of the team finalized their own experiments. The deployment was postponed to this morning as different minor bugs prevented us to do it yesterday evening.The GreenHab officer harvested the Spirulina that was growing since more than a week and started to dry it. Part of the crew left at 5 pm for an EVA to Grey Moon. Tonight we will all observe the stars, enjoying the clear sky that the Utah offers.

Look Ahead Plan: We planned two EVAs for tomorrow as it is our last day on Mars.

Part of the crew will go to North Ridge in the morning. The rest of the team will end their experiment and start writing the summary report. We will get back the antennas and the cubelanders, in the morning during an engineering EVA. In the afternoon, we will do a second EVA, again to North Ridge. The rest of the afternoon will be devoted to packing our cases.

Anomalies in work: No issues.

Weather: Sunny, almost no wind

Crew Physical Status: Good

EVA: The morning EVA along Hab Ridge Road was shortened by the presence of steam in our helmets due to the low temperature at this time of the day. We didn’t manage to find the sinkhole but we had nevertheless the occasion to take nice pictures with the morning light. The second EVA to Grey moon started at 5 pm. The EVA team saw breathtaking panoramas.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Request, EVA Report, Daily Photo Report

Support Requested: None

Operations Report – May 2nd

Crew 211 Operations Report 02-MAY-2019

SOL: 11

Name of person filing report: Julien Amalaberque

Non-nominal systems: The pump from the loft

Notes on non-nominal systems: * Loft pump: Still some weird noises but it holds. The leak is still present, nothing changed with that.

Generator (hours run last night): 11:00

On (last night): 8:00 pm

Off (this morning): 7:00 am

Solar— SOC 79 % at 19:30 pm

Diesel Reading – 100 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~190 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 65 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 15 gallons

Toilet tank emptied: no

Deimos rover used: no (out of order)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 92.1h

Beginning charge: 100

Ending charge: 66

Currently charging: yes

Opportunity rover used: no

Hours: 58h

Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used: yes

Hours: 93.3h

Beginning charge: 100

Ending charge: 71

Currently charging: yes

Notes on rovers: Nothing to report

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Some chards and carrots were harvested

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: As we started the food inventory, we noted that a large majority of the food is expired.

Questions, concerns and requests to Mission Support:

Copyright © The Mars Society. All rights reserved. | Main Site